OpenStudy (anonymous):

Solve for n: 8(n - 3) + 4n = 12n + 3

OpenStudy (shamil98):

Distribute the 8 to (n-3) first.

OpenStudy (anonymous):

8n-24

OpenStudy (shamil98):

Yep. 8n -24 +4n = 12n + 3 There are two terms you need to combine first then you can solve for n. 8n +4n + 24 = 12n + 3

OpenStudy (shamil98):

-24*

OpenStudy (anonymous):

12n-24 =12n+3 N=27?

OpenStudy (shamil98):

12n - 24 = 12n + 3 -12 n -12n -24 does not equal 3 so no solution on this one :P

OpenStudy (ranga):

^^^ correct. This problem does not have a solution.

OpenStudy (anonymous):

oh lol, thanks :)